OpenStudy (ammarah):

Find all real zeroes of the function. g(x)=2x^3+4x^2-2x-4

ganeshie8 (ganeshie8):

set it equal to 0 and factor out gcf if any

OpenStudy (ammarah):

1,2,4,8 then what?

ganeshie8 (ganeshie8):

wats 1,2,4,8 ?

ganeshie8 (ganeshie8):

to find zeroes, you need to solve below :- 2x^3+4x^2-2x-4 = 0

OpenStudy (ammarah):

i know but how do i soolve it?

ganeshie8 (ganeshie8):

factor 2 first

ganeshie8 (ganeshie8):

\(2x^3+4x^2-2x-4 = 0\) \(2(x^3+2x^2-x-2) = 0\) \(x^3+2x^2-x-2 = 0\)

ganeshie8 (ganeshie8):

Look at this cubic, by inspection 1 is a zero of this equation (why ? )

OpenStudy (ammarah):

so now i use synthetic division

ganeshie8 (ganeshie8):

Yup !

OpenStudy (ammarah):

okk so i used it and i got a zero by using 1 but then i dont know how to find thezeroes

ganeshie8 (ganeshie8):

wats the depressed equation you got after synthetic division ?

OpenStudy (ammarah):